a car starts from rest and attains a speed of 4m/s in 8s

a . calculate the acceleration of the car

b . calculate the distance it has covered during this time

c . what velocity would it attain at the end of 10s


1
Expert's answer
2020-11-19T06:01:31-0500

a) We can calculate the acceleration of the car from the kinematic equation:


"v=v_0+at,""a=\dfrac{v-v_0}{t}=\dfrac{4\ \dfrac{m}{s}-0\ \dfrac{m}{s}}{8\ s}=0.5\ \dfrac{m}{s^2}."

b) We can calculate the distance that the car has covered during this time from the another kinematic equation:


"d=v_0t+\dfrac{1}{2}at^2,""d=0\ \dfrac{m}{s}\cdot 8\ s+\dfrac{1}{2}\cdot 0.5\ \dfrac{m}{s^2}\cdot (8\ s)^2=16\ m."

c) We can calculate the the velocity that the car would attain at the end of 10 s from

the kinematic equation:


"v=v_0+at,""v(10\ s)=0\ \dfrac{m}{s}+0.5\ \dfrac{m}{s^2}\cdot 10\ s=5\ \dfrac{m}{s}."

Answer:

a) "a=0.5\ \dfrac{m}{s^2}."

b) "d=16\ m."

c) "v(10\ s)=5\ \dfrac{m}{s}."